# Res-VMamba: Fine-Grained Food Category Visual Classification Using Selective State Space Models with Deep Residual Learning

### v1

@@ 4.3 Implementation Details
Adhering to the protocol delineated in Liu et al. (2024a), Res-VMamba embarks on a comprehensive training regimen on CNFOOD-241, the backbone uses the VMamba-S, extending over 150 epochs with an initial warmup period covering the initial 20 epochs, and leverages a batch size of 128. The training schema integrates the AdamW optimizer, with the beta parameters set at $(0.9,0.999)$, and momentum fixed at 0.9. A cosine decay schedule modulates the learning rate, commencing with an initial learning rate of $1\times 10^{-3}$ and a weight decay parameter of 0.05. Augmenting the training are methodologies such as label smoothing at 0.1 and the implementation of an exponential moving average (EMA). Subsequent to these specified techniques, no further training strategies are deployed. The VMamba-S did transfer training with pretrained weight get from Liu et al. (2024a) on CNFOOD-241 with a batch size of 32, else training strategies is the same as default VMamba. The CMAL-Net are trained from ResNet-50 pretrained weight from pytorch based on its original setting on github, the others models are trained with pretrained weights on ImageNet-1K loaded from Huggingface by timm module with an initial learning rate of $1\times 10^{-4}$ and AdamW optimizer.
@@ 4.5 Ablation Study on Res-VMamba
Model
Year
Use PW.?
Top-1 Val. Acc.
Top-5 Val. Acc.
Top-1 Test Acc.
Top-5 Test Acc.
VGG16 Simonyan and Zisserman (2015)
2014
Y
66.98
90.10
65.06
89.60
ViT-B Dosovitskiy et al. (2021)
2020
Y
73.14
92.06
71.58
91.62
ResNet101 He et al. (2016)
2015
Y
74.42
93.62
72.59
93.16
DenseNet121 Huang et al. (2017)
2016
Y
76.46
94.57
74.77
94.29
Inceptionv4 Szegedy et al. (2016)
2016
Y
77.30
94.28
75.70
93.89
PRENet Min et al. (2023)
2017
Y*
77.47
94.86
76.02
94.61
SEnet154 Hu et al. (2018)
2017
Y
77.47
94.86
76.02
94.61
RepViT Wang et al. (2023)
2023
Y
78.08
95.41
76.86
95.02
ConvNeXT-B Liu et al. (2022)
2022
Y
78.30
94.36
76.76
93.90
EfficientNet-B6 Mingxing Tan (2019)
2019
Y
80.10
94.64
78.48
94.22
CMAL-Net Liu et al. (2023)
2023
${}^{\dagger}$
80.16
95.99
78.56
95.40
VMamba-S Liu et al. (2024a) 2024 N 79.17 95.64 77.73 95.24
@@
VMamba-S 2024 ${}^{\ddagger}$ 82.15 96.91 80.58 96.71
Table 3: Comparison of our approach (Res-Vmamba) to other baselines on CNFOOD-241 with our split method. PW., Val., Acc. denotes pretrained weight, validation, accuracy, respectively. *PRENet’s pretrained weight got from official github is trained on Food2K dataset. ${}^{\dagger}$ CMAL-Net used the ResNet-50 pretrained weight from pytorch, and other models’ pretrained weights got from timm module loaded from Huggingface trained on ImageNet-1K dataset. ${}^{\ddagger}$amba’s pretrained weight got from official github is trained on ImageNet-1K dataset.

### v2

@@ 4.1.2 Unbalanced Number of Images cross Categories on CNFOOD-241
Dataset Max. H Min. H Max. W Min. W Mean $\pm$ Std. H×$\times$×W
UEC Food100 800 71 800 80 358.25$\pm$139.46$\times$457.40$\pm$180.58
UEC Food256 800 71 800 80 406.99$\pm$117.16$\times$492.98$\pm$136.71
ETH Food-101 512 122 512 193 475.37$\pm$65.31$\times$495.79$\pm$45.67
UPMC Food-101 2960 120 2000 120 459.13$\pm$225.54$\times$559.80$\pm$268.67
Geo-Dish -* -* -* -* -*
UNICT-FD889 240 240 320 320 240$\pm$0.00$\times$320$\pm$0.00
Vireo Food-172 256 256 256 256 256$\pm$0.00$\times$256$\pm$0.00
Food11 6144 207 9542 220 496.01$\pm$180.38$\times$531.55$\pm$250.92
UNICT-FD1200 240 240 320 320 240$\pm$0.00$\times$320$\pm$0.00
ChineseFoodNet -* -* -* -* -*
Vegfru 6016 45 7360 48 309.73$\pm$272.46$\times$374.52$\pm$338.21
Sushi-50 4752 232 6570 285 767.20$\pm$584.50$\times$955.80$\pm$758.16
FoodX-251 2744 256 2733 256 287.48$\pm$75.45$\times$341.24$\pm$86.28
ISIA Food-200 8868 70 10030 78 707.28$\pm$561.74$\times$830.91$\pm$654.40
Taiwanese-Food-101 7360 61 6720 83 613.76$\pm$372.83$\times$771.91$\pm$455.90
ISIA Food-500 8868 100 10361 78 700.78$\pm$566.41$\times$822.08$\pm$667.43
Food2K 1489 75 1717 97 381.26$\pm$74.96$\times$ 515.89$\pm$102.72
CNFOOD-241 600 600 600 600 600$\pm$0.00$\times$600$\pm$0.00
Table 2: Image size statistics comparison of current open datasets of food recognition (pixels). Max., Min., Std. are maximum, minimum and standard deviation respectively. H, W denotes the height and width of image respectively. The CNFOOD-241 dataset has the largest balenced image size in all the food image dataset we were available. *The data download link on the Geo-Dish (http://isia.ict.ac.cn/dataset/Geolocation-food/) and ChineseFoodNet (https://goo.gl/kWNV8a) official pages are dead, so we can not get the dataset to analyze image properties.
@@ 4.3 Implementation Details
Adhering to the protocol delineated in [15], Res-VMamba embarks on a comprehensive training regimen on CNFOOD-241, the backbone uses the VMamba-S, extending over 150 epochs with an initial warmup period covering the initial 20 epochs, and leverages a batch size of 128. The training schema integrates the AdamW optimizer, with the beta parameters set at $(0.9,0.999)$, and momentum fixed at 0.9. A cosine decay schedule modulates the learning rate, commencing with an initial learning rate of $1\times 10^{-3}$ and a weight decay parameter of 0.05. Augmenting the training are methodologies such as label smoothing at 0.1 and the implementation of an exponential moving average (EMA). Subsequent to these specified techniques, no further training strategies are deployed. The VMamba-S did transfer training with pretrained weight get from [15] on CNFOOD-241 with a batch size of 32, else training strategies is the same as default VMamba. The CMAL-Net are trained from ResNet-50 pretrained weight from pytorch based on its original setting on github, the others models are trained with pretrained weights on ImageNet-1K loaded from Huggingface by timm module with an initial learning rate of $1\times 10^{-4}$ and AdamW optimizer.
@@ 4.5 Ablation Study on Res-VMamba
Model Year Use PW.? Top-1 Val. Acc. Top-5 Val. Acc. Top-1 Test Acc. Top-5 Test Acc.
VGG16 [68] 2014 Y 66.98 90.10 65.06 89.60
ViT-B [69] 2020 Y 73.14 92.06 71.58 91.62
ResNet101 [65] 2015 Y 74.42 93.62 72.59 93.16
DenseNet121 [70] 2016 Y 76.46 94.57 74.77 94.29
Inceptionv4 [71] 2016 Y 77.30 94.28 75.70 93.89
PRENet [41] 2017 Y* 77.47 94.86 76.02 94.61
SEnet154 [48] 2017 Y 77.47 94.86 76.02 94.61
RepViT [47] 2023 Y 78.08 95.41 76.86 95.02
ConvNeXT-B [45] 2022 Y 78.30 94.36 76.76 93.90
EfficientNet-B6 [44] 2019 Y 80.10 94.64 78.48 94.22
CMAL-Net [43] 2023 Y 80.16 95.99 78.56 95.40
VMamba-S [15] 2024 N 79.17 95.64 77.73 95.24
@@
VMamba-S 2024 Y 82.15 96.91 80.58 96.71
Table 4: Comparison of our approach (Res-Vmamba) to other baselines on CNFOOD-241 with our split method. PW., Val., Acc. denotes pretrained weight, validation, accuracy, respectively. *PRENet’s pretrained weight got from official github is trained on Food2K dataset. CMAL-Net used the ResNet-50 pretrained weight from pytorch, and other models’ pretrained weights got from timm module loaded from Huggingface trained on ImageNet-1K dataset. VMamba’s pretrained weight got from official github is trained on ImageNet-1K dataset.